# Decoding News Narratives: A Critical Analysis of Large Language Models in Framing Bias Detection

### v2

@@ Abstract
This work contributes to the expanding research on the applicability of LLMs in social sciences by examining the performance of GPT-3.5 Turbo, GPT-4, and Flan-T5 models in detecting framing bias in news headlines through zero-shot, few-shot, and explainable prompting methods. A key insight from our evaluation is the notable efficacy of explainable prompting in enhancing the reliability of these models, highlighting the importance of explainable settings for social science research on framing bias. GPT-4, in particular, demonstrated enhanced performance in few-shot scenarios when presented with a range of relevant, in-domain examples. FLAN-T5’s poor performance indicates that smaller models may require additional task-specific fine-tuning for identifying framing bias detection. Our study also found that models, particularly GPT-4, often misinterpret emotional language as an indicator of framing bias, underscoring the challenge of distinguishing between reporting genuine emotional expression and intentionally use framing bias in news headlines. We further evaluated the models on two subsets of headlines where the presence or absence of framing bias was either clear-cut or more contested, with the results suggesting that these models’ can be useful in flagging potential annotation inaccuracies within existing or new datasets. Finally, the study evaluates the models in real-world conditions ("in the wild"), moving beyond the initial dataset focused on U.S. Gun Violence, assessing the models’ performance on framed headlines covering a broad range of topics111All data created in this study will be made publicly available..
Decoding News Narratives: A Critical Analysis of Large Language Models in Framing Bias Detection
@@ 1 Introduction
In today’s digital age, the proliferation of news sources and the rapid dissemination of information highlight the essential need for unbiased reporting. Framing bias, which manipulates news content to potentially shift public perception, presents a formidable obstacle in maintaining a well-informed and unbiased public sphere Binotto and Bruno (2018). This form of bias not only influences how events are portrayed but also impacts public attitudes and policy decisions. For instance, the reporting on a government policy change might be emotionally charged in one headline as “Government’s heartless cutbacks leave thousands without essential services”, whereas another might offer a more neutral perspective: “Government announces reduction in funding for public services”.
@@
The concept of framing has received extensive scrutiny in the social sciences, identified as a technique that selectively highlights certain realities to shape audience perceptions and reactions Goffman (1974); Entman (1993). While its significant effect on public communication and media is well-documented, the analysis of framing has historically been challenged by its fragmented nature and reliance on manual, small-scale studies. This situation underscores the necessity for new methods capable of dissecting the complexities of framing in media comprehensively. Amid this landscape, with an overwhelming array of news sources and headlines, it becomes crucial to determine whether they convey biased messages, intentionally or inadvertently. This determination is not merely academic but essential for preserving the integrity of public discourse.
The introduction of large language models (LLMs) opens up novel pathways for identifying framing bias, especially in news headlines. The accessibility and advanced capabilities of models, particularly those within the GPT series, have garnered attention from professionals across various fields, including media analysis and political science, seeking to navigate the extensive corpus of available information. However, despite their impressive success across many applications, the reliability of these models in executing nuanced tasks like detecting framing bias remains an open question. In this study, we address this gap by providing an in-depth investigation into the strengths and shortcomings of state-of-the-art natural language processing (NLP) models in identifying framing bias. We perform a comprehensive examination of the performance of two top-tier NLP models, GPT-4 and GPT-3.5, alongside the open-source FLAN-T5 model (Wei et al., 2022), which has a reasonable performance across different NLP tasks.
Our findings reveal interesting patterns: while the pre-trained FLAN-T5 struggles with this task, the GPT models show promising results. Our results suggest that prompting the models to explain their decisions results in more reliable performances across different settings. A standout discovery is the enhanced reliability of predictions when models are prompted to explain their reasoning, a technique that also narrows the variability in outcomes across different example sets. To test the real-world applicability of these models, apart from a standard dataset that is annotated by experts, we automatically collect a new evaluation set containing framed headlines from different domains. The GPT models also accurately identifying the majority of framed titles in this dataset.
Our analysis suggests that while the examined GPT models achieve high scores on the examined evaluation sets, there are still areas that they struggle. As an example, GPT-4 often mistook emotional language for framing bias. This insight points to the need for developing more nuanced test sets that can present more complex challenges for evaluation. Additionally, our experiments suggest a potential use case for these models beyond mere bias detection. When GPT models, particularly in a more reliable prompting setting, converge on a prediction that deviates from the established gold standard, it may hint at underlying annotation inaccuracies within the dataset. Our findings suggest that a future direction involves constructing more challenging datasets that mirror the real-world complexities of framing bias, as well as enhancing the performance of smaller and open-source models on this challenging task.
@@ 2.1 Automatic Framing Detection
Various NLP methodologies have been used to automate framing bias detection. Early studies in framing detection primarily utilized Topic Modeling, Structural Topic Modeling, and Hierarchical Topic Modeling to uncover themes and topics within large datasets DiMaggio et al. (2013); Nguyen et al. (2015); Gilardi et al. (2021). These methods, while effective in identifying “what” is being discussed, often fall short in revealing “how” information is framed. Latent Dirichlet Allocation (LDA) Topic Modeling, for instance, served as a starting point for creating lists of frames deductively in tools like the one presented by Bhatia et al. (2021) for computational framing analysis. However, as noted by Ali and Hassan (2022), the emphasis in such approaches remains on detecting topics rather than the nuanced framing of those topics. The focus on topics also derives from the connections between agenda setting and framing strategies in computational social sciences, with studies analysing these two phenomena together Field et al. (2018). Further analyses have included pragmatics cues, examining how specific word choices, like the use of “again” in “Again, Dozens of Refugees Drowned”, subtly influence reader perception Yu (2022). This shift towards granular analysis is complemented by advanced models, including Neural Network and deep learning techniques, which offer refined tools for detecting framing nuances Burscher et al. (2016); Card et al. (2015); Liu et al. (2019); Mendelsohn et al. (2021). Tourni et al. (2021) demonstrated that combining transformers models for processing news headlines with residual network models to process news lead images could lead to accurate framing bias detection. Similarly, Naderi and Hirst (2017) explored the use of various deep neural networks, such as LSTMs, BiLSTMs, and GRUs, for frame prediction at the sentence level using the Media Frame Corpus (MFC) Card et al. (2015). Building on this foundation, Liu et al. (2019) and Akyürek et al. (2020) fine-tuned BERT Devlin et al. (2019) to predict frames in news headlines. Their work resulted in the creation of the Gun Violence Frame Corpus (GVFC), a benchmark dataset for framing analysis which will be further discussed in section 2.1.1. More recently, in a novel approach to automatic framing detection, Lai et al. (2022) introduced an unsupervised learning method leveraging Wikipedia’s category system to identify frames within news articles, using the GVFC to evaluate this method.
2.1.1 Datasets
@@
MFC (Card et al., 2015) is a collection of annotated U.S. newspaper articles on topics like immigration, smoking, and same-sex marriage, analyzed for framing. Utilising the Policy Frames Codebook (PFC) by Boydstun et al. (2014), the MFC adopts 14 frame dimensions such as “security and defense” and “cultural identity” for categorising policy discourse. Despite achieving an inter-coder reliability (ICR) of 0.60, critiques, particularly from Ali and Hassan (2022), argue that the PFC’s broad dimensions conflate topics with frames, potentially missing nuanced strategic framing. Moreover, MFC categorises content into wide-ranging dimensions that might not always precisely capture the specific framing intended by a news headline. This categorisation can make it difficult to directly identify whether and how a headline is framed without a deeper, nuanced analysis.
@@
Another significant dataset in the field of framing analysis is the Gun Violence Frame Corpus (GVFC), introduced by Liu et al. (2019). This dataset concentrates on the issue of Gun Violence in the U.S. The creation process began with defining nine distinct “frames” related to the topic, drawing from existing literature and a preliminary data analysis. A specialised codebook was then developed, serving as a training tool for annotators along with annotation guidelines. GVFC is made of 2990 news headlines, with 1,300 headlines specific to the issue of Gun Violence in the United States. All the headlines are coded to have a primary frame, while only 319 have 2 frames. For instance, the headline “It’s Time to Hand the Mic to Gun Owners” is annotated with “Public opinion” as the first frame and “2nd Amendment” as the second frame. Similarly, “Trevor Noah: The Second Amendment Is Not Intended for Black People” is annotated with “2nd Amendment” and “Race/Ethnicity” frames (Liu et al., 2019, p. 507).
@@
Expanding the scope of framing analysis to non-English content, Akyürek et al. (2020) introduced a multilingual extension of the Gun Violence Frame Corpus, which encompasses news headlines in German, Turkish, and Arabic, focusing on U.S. gun violence. This extension involved training two native speakers per language to annotate headlines—350 in German, 200 in Turkish, and 210 in Arabic.
Piskorski et al. (2023b) presented an annotated dataset made of articles spanning though nine languages: English, French, German, Georgian, Greek, Italian, Polish, Russian, and Spanish. This dataset addresses a variety of topics including the COVID-19 pandemic, abortion-related legislation, migration, Russo-Ukrainian war, and various parliamentary elections. The annotation process made use of the PFC codebook, using the 15 dimensions as frames Piskorski et al. (2023a).
While our study focuses on the English language, exploring framing bias detection in less resourced languages presents a promising future direction. By establishing a foundation in English, we lay the groundwork for future research to investigate the strengths and shortcomings of framing bias detection in languages beyond English.
@@ 2.2 Evaluating LLMs in Social Science
Applying Large Language Models to social science tasks, such as evaluating sociability Choi et al. (2023), morality Abdulhai et al. (2023), and controversial issues and bias Sun et al. (2023), has received increasing interest, showcasing a wide range of strengths and limitations unique to each task. This diversity stems from the specific challenges and nuances of social phenomena. Although LLMs excel in generating and understanding human-like text, the complex requirements of social science tasks necessitate a detailed, task-specific examination of their performance and reliability.
In this work, we contribute to the expanding research on the applicability of LLMs in social sciences by specifically investigating their reliability in detecting framing bias. This exploration not only aims to identify the strengths and shortcomings of state-of-the-art models in a crucial area of framing analysis but also to extend our comprehension of how these models can be optimised for the nuanced task of framing detection.
@@ 3.1 Data
For our evaluation, we selected the Gun Violence Framing Corpus (GVFC) dataset Liu et al. (2019), motivated by its comprehensive coverage of U.S. Gun Violence framing through 2900 annotated newspaper headlines as well as the high ICR met in the annotation process.222In our experiments, we have excluded the headlines that are not relevant to Gun Violence. Furthermore, we have excluded 22 relevant headlines in order to use them for Few-Shot in domain prompting, leaving 2594 relevant headlines for our analysis. The dataset’s annotations identify whether each headline reflects any of nine critical aspects of gun violence framing: gun rights, gun control, politics, mental health, public/school safety, race/ethnicity, public opinion, social/cultural issues, and economic consequences. In our study, headlines tagged with any of these framing aspects were categorised as framed, while all others were classified as not framed.
@@ 3.2 Models
To assess the capabilities of state-of-the-art NLP models in identifying framing bias, our study focuses on two primary categories of models: the GPT-4333GPT-4-0613 OpenAI (2023) and GPT-3.5-Turbo444GPT-3.5-turbo-0613 Ye et al. (2023), known for their state-of-the-art performance across a wide range of NLP benchmarks, and the open-source FLAN-T5 Wei et al. (2022), which, despite its accessibility, demonstrates significant efficacy in various tasks Chung et al. (2022). Additionally, we examined various sizes of the FLAN-T5 model, i.e., small (77M parameters), base (248M parameters), and large (783M parameters), to explore how model scale influences bias detection capabilities.
@@
Our evaluation methodology employs a uniform series of prompts applied across all models under three distinct experimental conditions: (1) a zero-shot setting, probing the models’ inherent knowledge on framing bias detection; (2) a few-shot scenario, wherein models are provided with a small set of examples to inform their bias detection process; and (3) an explainability-focused approach, where models are asked to explain their rationale behind their decisions aside from label prediction.
@@
Our evaluation framework is designed to thoroughly evaluate the models’ adeptness at navigating tasks without being finetuned on framing bias datasets. This aspect is particularly important within the domain of social science research, where specialised training datasets are often scarce.
3.2.1 Zero-Shot Prompting
@@
In zero-shot settings, we deploy two distinct approaches to evaluate the models’ effectiveness in detecting framing bias. The first approach involves presenting the models with a straightforward task: determining if a headline is framed, without any additional context or examples.555This involves using the “Decide whether this claim is framed:” prompt for GPT models, and “Is this claim framed? OPTIONS Yes — No” for FLAN-T5. We chose the FLAN-T5 model’s prompt to ensure it aligns with the instructions that this model encountered during pretraining.
@@
In the enhanced second zero-shot setting, we further add a specific definition to the input prompt to guide the task of framing detection. We use the “a communication strategy often used in journalism and political language, where certain aspects of an issue are highlighted while others are minimised or ignored, thereby promoting a particular interpretation of that issue” definition for this purpose.666This definition is grounded on the existing definitions of framing by Goffman (1974); Entman (1993, 2007). The model is then tasked with deciding if a given claim is framed based on this definition. This additional guidance aims to refine the models’ analytical lens, providing a clearer framework for identifying and evaluating framing within texts.
3.2.2 Few-Shot Prompting

### v3

# Decoding News Narratives: A Critical Analysis of Large Language Models in Framing Detection

@@ Abstract
Previous studies on framing have relied on manual analysis or fine-tuning models with limited annotated datasets. However, pre-trained models, with their diverse training backgrounds, offer a promising alternative. This paper presents a comprehensive analysis of GPT-4, GPT-3.5 Turbo, and FLAN-T5 models in detecting framing in news headlines. We evaluated these models in various scenarios: zero-shot, few-shot with in-domain examples, cross-domain examples, and settings where models explain their predictions. Our results show that explainable predictions lead to more reliable outcomes. GPT-4 performed exceptionally well in few-shot settings but often misinterpreted emotional language as framing, highlighting a significant challenge. Additionally, the results suggest that consistent predictions across multiple models could help identify potential annotation inaccuracies in datasets. Finally, we propose a new small dataset for real-world evaluation on headlines from a diverse set of topics.111All data created in this study will be made publicly available.
Decoding News Narratives: A Critical Analysis of Large Language Models in Framing Detection
@@ 1 Introduction
In today’s digital age, the spread of news sources and the rapid dissemination of information highlight the essential need for unbiased reporting. Framing is a communication strategy used to manipulate news content in order to potentially shift public perception to follow a certain agenda, and is an obstacle in maintaining a well-informed and unbiased public audience Binotto and Bruno (2018). Framing not only influences how events are perceived and remembered but can also impact overall attitudes and policy decisions, since it helps building a potentially biased narrative around an event. For instance, the reporting on a government policy change might be framed in one headline as “Government’s heartless cutbacks leave thousands without essential services”, whereas another might offer a more neutral perspective: “Government announces reduction in funding for public services”.
@@
The concept of framing has received extensive analysis in social science, identified as a technique that selectively highlights certain realities to shape audience perceptions and reactions Goffman (1974); Entman (1993). Despite its well-documented impact on public communication and media, analysing framing has traditionally been limited by manual, small-scale studies Baumer et al. (2015). This limitation highlights the urgent need for new methods capable of thoroughly analysing the complexities of framing in media. In today’s context, with an overwhelming array of news sources and headlines, it is crucial to determine whether they convey framed messages. This task is not merely academic but essential for preserving the integrity of public discourse.
The introduction of large language models (LLMs) provides a promising alternative for identifying framing bias in news headlines. LLMs, already trained on vast corpus of text data, can be readily adapted or tested for similar tasks using prompts or a few examples. Despite their success in many areas, their reliability in handling nuanced tasks like detecting framing remains an open question.
This study aims to address this gap by providing an in-depth investigation into the strengths and shortcomings of prominent instruction-following pre-trained language models in identifying framing in news headlines. We perform a comprehensive examination of the performances of GPT-4, GPT-3.5 Turbo, and open source FLAN-T5 model. We chose FLAN-T5 due to its robust performance across various NLP tasks (Wei et al., 2022) and its status as one of the first models designed to follow instructions or prompts effectively.
Our results reveal that, while out-of-the-box FLAN-T5 models struggle with framing detection, the GPT models show promising results. Notably, prompting the models to explain their decisions results in more reliable performance across different settings.
However, our analysis indicates that while the GPT models perform well on the evaluation sets, they still face some challenges. For instance, GPT-4 often mistook emotional language for framing. This insight underscores the need for more nuanced framing test sets that present complex challenges for evaluation.
Additionally, our experiments suggest a potential use case for these models beyond mere framing detection. In fact, by dividing the dataset into clear-cut and contested subsets based on the original annotations, we observed that when GPT models, particularly in a more reliable prompting setting, agree in deviating from the established gold standard, it may indicate underlying annotation inaccuracies.
To assess the real-world applicability of these models across various domains, we collected a new evaluation set containing framed headlines from different topics, achieving an accuracy of 75.16% for GPT-3.5 Turbo and 80.25% for GPT-4.
Finally, our study shows that, when fine-tuned, FLAN-T5 outperforms both GPT models on in-domain data (GVFC) with an accuracy of 78.83% compared to GPT-4’s 66.92%. However, this advantage diminishes with out-of-domain evaluation in real-world settings, where its accuracy drops to 68.78% compared to GPT-4’s 80.25%. This emphasises the effectiveness of prompt-based techniques without fine-tuning for analysing the presence of framing in real world scenarios.
@@ 2.1 Automatic Framing Detection
Various NLP methodologies have been used to automate framing detection. Early studies in framing detection primarily utilised Topic Modeling, Structural Topic Modeling, and Hierarchical Topic Modeling to uncover themes and topics within large datasets DiMaggio et al. (2013); Nguyen et al. (2015); Gilardi et al. (2021). These methods, while effective in identifying “what” is being discussed, often fall short in revealing “how” information is framed.
Latent Dirichlet Allocation (LDA) Topic Modeling, for instance, served as a starting point for creating lists of frames deductively in tools like the one presented by Bhatia et al. (2021) for computational framing analysis. However, as noted by Ali and Hassan (2022), the emphasis in such approaches remains on detecting topics rather than the nuanced framing of those events. The focus on topics could also derive from the connections between agenda setting and framing strategies in computational social sciences, with studies analysing these two phenomena together Field et al. (2018).
Further analyses have included pragmatics cues, examining how specific word choices, like the use of “again” in “Again, Dozens of Refugees Drowned”, subtly influence reader perception Yu (2022). This shift towards granular analysis is complemented by advanced models, including Neural Network and deep learning techniques, which offer refined tools for detecting framing nuances Burscher et al. (2016); Card et al. (2015); Liu et al. (2019); Mendelsohn et al. (2021). Tourni et al. (2021) demonstrated that combining transformer models for processing news headlines with residual network models to process news lead images could improve the accuracy of framing detection. Similarly, Naderi and Hirst (2017) explored the use of various neural networks, such as LSTMs, BiLSTMs, and GRUs, for frame prediction at the sentence level using the Media Frame Corpus (MFC) Card et al. (2015).
Building on this foundation, Liu et al. (2019) and Akyürek et al. (2020) fine-tuned BERT Devlin et al. (2019) to predict frames in news headlines. Their work resulted in the creation of the Gun Violence Frame Corpus (GVFC), a benchmark dataset for framing analysis which will be further discussed in section 2.2. More recently, in a novel approach to automatic framing detection, Lai et al. (2022) introduced an unsupervised learning method leveraging Wikipedia’s category system to identify frames within news articles, using the GVFC to evaluate this method.
2.2 Datasets
@@
MFC (Card et al., 2015) is a collection of annotated U.S. newspaper articles on topics like immigration, smoking, and same-sex marriage, analysed for framing. Utilising the Policy Frames Codebook (PFC) by Boydstun et al. (2014), the MFC adopts 14 frame dimensions such as “security and defense” and “cultural identity” for categorising policy discourse. Despite achieving an inter-coder reliability (ICR) of 0.60, critiques, particularly from Ali and Hassan (2022), argue that the PFC’s broad dimensions conflate topics with frames, potentially missing nuanced strategic framing.
Moreover, MFC categorises content into wide-ranging dimensions (i.e., politics, economic, etc.) that might not always precisely capture the specific framing intended by a news headline Ali and Hassan (2022). This categorisation can make it difficult to directly identify whether and how a headline is framed without a deeper, nuanced analysis. For example, the headline “Where the jobs are” has been annotated using the framing dimension “Economic”, while the following portion of text “Critics of illegal immigration can make many cogent arguments to support the position that the U.S. Congress and the Colorado legislature must develop effective and well-enforced immigration policies that will restrict the number of people who migrate here legally and illegally.” was annotated using the framing dimension “Policy prescription”.
@@
Another significant dataset in the field of framing analysis is the GVFC, introduced by Liu et al. (2019). This dataset concentrates on the issue of Gun Violence in the U.S. The creation process began with defining nine distinct “frames” related to the topic, drawing from existing literature and a preliminary data analysis. A specialised codebook was then developed, serving as a training tool for annotators along with annotation guidelines.
GVFC is made of 2990 news headlines, with 2616 headlines specific to the issue of Gun Violence in the United States. All the in-domain headlines are coded to have a primary frame, while only 319 have two frames. For instance, the headline “It’s Time to Hand the Mic to Gun Owners” is annotated with “Public opinion” as the first frame and “2nd Amendment” as the second frame. Similarly, “Trevor Noah: The Second Amendment Is Not Intended for Black People” is annotated with “2nd Amendment” and “Race/Ethnicity” frames (Liu et al., 2019).
@@
Expanding the scope of framing analysis to non-English content, Akyürek et al. (2020) introduced a multilingual extension of the Gun Violence Frame Corpus, which encompasses news headlines in German, Turkish, and Arabic, focusing on U.S. gun violence. This extension involved training two native speakers per language to annotate headlines - 350 in German, 200 in Turkish, and 210 in Arabic.
Piskorski et al. (2023b) presented an annotated dataset made of articles spanning through nine languages: English, French, German, Georgian, Greek, Italian, Polish, Russian, and Spanish. This dataset addresses a variety of topics including the COVID-19 pandemic, abortion-related legislation, migration, Russo-Ukrainian war, and various parliamentary elections. The annotation process made use of the PFC codebook, using the 15 dimensions as frames Piskorski et al. (2023a).
@@ 2.3 Evaluating LLMs in Social Science
Applying Large Language Models to social science tasks, such as evaluating sociability Choi et al. (2023), morality Abdulhai et al. (2023), and controversial issues and bias Sun et al. (2023), has received increasing interest, showcasing a wide range of strengths and limitations of the examined language models unique to each task. This diversity stems from the specific challenges and nuances of social phenomena. Although LLMs excel in generating and understanding human-like text, the complex requirements of social science tasks necessitate a detailed, task-specific examination of their performance and reliability.
In this work, we contribute to the expanding research on the applicability of LLMs in social sciences by specifically investigating their reliability in detecting framing.
@@ 3.1 Data
For our evaluation, we selected GVFC Liu et al. (2019), motivated by its comprehensive coverage of U.S. Gun Violence framing as well as the high ICR met in the annotation process. In our experiments, we have excluded the headlines that are not relevant to Gun Violence and hence, are not annotated with framing information. The dataset’s annotations identify whether each headline reflects any of nine critical aspects of gun violence framing: gun rights, gun control, politics, mental health, public/school safety, race/ethnicity, public opinion, social/cultural issues, and economic consequences. We consider headlines tagged with any of the above categories as framed, while all others as not framed. Furthermore, we have excluded 22 relevant headlines in order to use them for Few-Shot in domain prompting, leaving 2594 relevant headlines for our analysis. Of these, 1293 are framed and 1301 are not framed. Hence, the majority label is Not Framed and the majority baseline is 50.15%.
@@ 3.2 Models
To assess the capabilities of two of the leading closed-source models in identifying framing, our study focuses on the GPT-4222GPT-4-0613. Temperature parameter set to 0. OpenAI (2023) and GPT-3.5-Turbo333GPT-3.5-turbo-0613. Temperature parameter set to 0. Ye et al. (2023), known for their state-of-the-art performance across a wide range of NLP benchmarks. We chose GPT-4 and GPT-3.5 Turbo due to their advanced capabilities and ease of use. While they are closed-source with a paid API, they are easily accessible for social scientists, eliminating the need for setting up libraries and running models with high memory and expertise requirements. This accessibility makes them practical tools for researchers who may not have extensive computational resources.
@@
Additionally, we chose the open-source FLAN-T5 Wei et al. (2022), which, despite its accessibility, demonstrates significant efficacy in various tasks Chung et al. (2022). We examined various sizes of the FLAN-T5 model, i.e., small (77M parameters), base (248M parameters), and large (783M parameters), to explore how model scale influences framing detection capabilities.
@@
Our evaluation methodology employs a uniform series of prompts applied across all models under three distinct experimental conditions: (1) a zero-shot setting, probing the models’ inherent knowledge on framing detection; (2) a few-shot scenario, wherein models are provided with a small set of examples to inform their framing detection process; and (3) an explainability-focused approach, where models are asked to explain their rationale behind their decisions aside from label prediction.
@@
Our evaluation framework is designed to thoroughly evaluate the models’ adeptness at navigating tasks without being fine-tuned on framing datasets. This aspect is particularly important within the domain of social science research, where specialised training datasets are often scarce.
3.3 Zero-Shot Prompting
In zero-shot settings, we deploy two distinct approaches to evaluate the models’ effectiveness in detecting framing.
@@ 3.3 Zero-Shot Prompting
The first approach involves presenting the models with a straightforward task: determining if a headline is framed, without any additional context or examples.444GPT prompt: “Decide whether this claim is framed”.
FLAN-T5 prompt:“Is this claim framed? OPTIONS Yes — No”. We chose the FLAN-T5 model’s prompt to ensure it aligns with the instructions that this model encountered during pre-training.
@@
In the enhanced second zero-shot setting, we further add a specific definition to the input prompt to guide the task of framing detection. We use the “a communication strategy often used in journalism and political language, where certain aspects of an issue are highlighted while others are minimised or ignored, thereby promoting a particular interpretation of that issue” definition for this purpose.555This definition is grounded on the existing definitions of framing by Goffman (1974); Entman (1993, 2007). The model is then tasked with deciding if a given claim is framed based on this definition. This additional guidance aims to refine the models’ analytical lens, providing a clearer framework for identifying and evaluating framing within texts.
3.4 Few-Shot Prompting